x+x-5 greater than or equal to < 33
1
Expert's answer
2011-10-25T10:23:59-0400
x+x-5&gt;=33



=&gt; 2x-5&gt;=33

let's add 5 to each part of the inequality:

2x-5+5&gt;=33+5

or& 2x&gt;=38

dividing each part of the inequality by 2 (sign of the inequality won't change cause 2&gt;0) we obtain

x&gt;=19 - answer of the question